# Automatically Labeling \$200B Life-Saving Datasets: A Large Clinical Trial Outcome Benchmark

### v1

@@ Abstract
The global cost of drug discovery and development exceeds \$200 billion annually. The main results of drug discovery and development are the outcomes of clinical trials, which directly influence the regulatory approval of new drug candidates and ultimately affect patient outcomes. Despite their significance, large-scale, high-quality clinical trial outcome data are not readily available to the public. Suppose a large clinical trial outcome dataset is provided; machine learning researchers can potentially develop accurate prediction models using past trials and outcome labels, which could help prioritize and optimize therapeutic programs, ultimately benefiting patients. This paper introduces Clinical Trial Outcome (CTO) dataset, the largest trial outcome dataset with around $479$K clinical trials, aggregating outcomes from multiple sources of weakly supervised labels, minimizing the noise from individual sources, and eliminating the need for human annotation. These sources include large language model (LLM) decisions on trial-related documents, news headline sentiments, stock prices of trial sponsors, trial linkages across phases, and other signals such as patient dropout rates and adverse events. CTO’s labels show unprecedented agreement with supervised clinical trial outcome labels from test split of the supervised TOP dataset [11], with a $91$ F1.
1 Introduction
@@
A clinical trial is an indispensable step toward developing a new drug, involving human participants to test the drug’s efficacy and safety for treating target diseases. In 2022, drug discovery and development spending reached 244 billion dollars globally [38]. Among which, the clinical trial market reached \$44.3 billion in 2020 and is expected to grow to \$69.3 billion by 2028 [33]. Low efficacy, safety issues, and poor trial protocol design can lead to trial failures [10, 37, 25]. Eroom’s Law111reverse of “Moore’s Law” shows that the number of new FDA-approved drugs per billion US dollars of R&D spending has halved approximately every nine years since 1950, even with inflation adjustment [35]. Given these challenges, predicting trial outcomes in silico—using computational methods—could significantly enhance drug discovery efficiency.
Surprising challenge: Despite the significant effort and resources invested in clinical trials, it is surprising that trial outcomes are not readily available for all trials. While some trial results have been published, many do not have publications available, and connecting the trials across different phases is also challenging. This lack of high-quality trial outcome labels presents a major obstacle in creating predictive models related to trial outcomes, which could potentially optimize the drug development process and improve patient outcomes. Furthermore, the FDA does not release the clinical trial ID (NCTID) in documents of approved drug applications.
Public data sources, such as the ClinicalTrials.gov database with more than 400,000 historical trials [34, 46, 47], provide vital information for identifying trial outcome labels. Other valuable resources include the Food and Drug Administration (FDA) National Drug Code (NDC) directory [9], which offers a comprehensive set of drug approvals and their codes, and the DrugBank [44, 23] database, which contains biochemical descriptions of many drugs and their indications. However, these resources are not connected, and the lack of direct links between clinical trials, drug application processes, and different phases of drug interventions makes it difficult to obtain clear trial outcome labels. The absence of a centralized, easily accessible database that consolidates clinical trial outcomes, drug approvals, and intervention phases poses a significant challenge for researchers and drug developers. This fragmented landscape of information hinders the development of accurate predictive models and can slow down the drug discovery and development process.
@@
Trial Outcome Definitions Clinical trial outcomes are multifaceted and have diverse implications. These outcomes can involve meeting the primary endpoint as defined in the study, advancing to the next phase of the trial, obtaining regulatory approval, impacting the financial outcome for the sponsor (either positively or negatively), and influencing patient outcomes such as adverse effects and trial dropouts. Our paper follows the previous conventions [11, 41, 10, 24, 1] and defines the trial outcome as a binary indicator, showing whether the trial achieves its primary endpoints and can progress to the next stage of drug development. For Phase 1 and 2 trials, success may mean moving to the next phase, such as from Phase 1 to Phase 2, and from Phase 2 to Phase 3. In Phase 3, success is measured by regulatory approval.
@@
Recently, initial efforts have been made to forecast specific aspects of clinical trials to enhance their outcomes. These efforts include employing electroencephalographic (EEG) measures to predict the effects of antidepressant treatments [30], optimizing drug toxicity predictions based on drug and target properties [16], and using phase II trial results to anticipate phase III trial outcomes [29]. Additionally, there is a growing interest in developing comprehensive methods for predicting trial outcomes. For example, predicting drug approvals for 15 disease groups by analyzing drug and clinical trial features using classical machine learning techniques [24], using multimodal drug structure and text information to predict outcomes based on a supervised set of data [11], an algorithm that computes the probability of technical success via asking experts a standardized questionnaire [43], and multimodal trial outcome prediction via omics, text, clinical trial design, and small molecule properties [1]. Despite these efforts, several limitations still impede the utility of existing trial outcome prediction models. Namely–the lack of transparency in the clinical trial labeling process. Fu et al. [11] is one example of a publicly available expert-curated clinical trial dataset. However, it is also quite limited because it only contains 17,538 human-labeled interventional small-molecule drug trials, out of around 400,000 total trials. To date, we are unaware of any other large-scale, publicly available, open-source (fully reproducible) effort to compute trial outcome labels.
We state our contributions as follows:
  • We propose CTO, the first large-scale, publicly available, open-sourced, and fully reproducible dataset of clinical trial outcomes derived from multiple sources of weakly supervised labels, including trial phase linkages, LLM interpretations of trial-related publications, news headline sentiments, stock prices of trial sponsors, and other trial metrics.
  • CTO demonstrates significant agreement with published trial outcome results, achieving promising results with a 94 F1 score on Phase 3 trials and 91 F1 score on all phase trials compared to humanly annotated clinical trial outcomes.
  • We provide all code [^2] https://github.com/chufangao/CTOD/ and data in a reproducible and easily extendable format, allowing for calculating predicted labels for new trials. Additionally, we aggregate trial-related data, such as ICD coding, drug mapping, and publications, facilitating secondary applications and benchmarking current outcome prediction models.
Figure 1: CTO integrates various data sources to generate labels for predicting clinical trial outcomes. Data sources include (1) News articles, from which trial embeddings and sentiment scores are derived; (2) Publicly available stock prices, used to compute price changes before and after trials; (3) PubMed abstracts, where LLMs are prompted to predict the trial outcome; and (4) trial phase linkage. Additional sources of outcomes include trial metrics such as the number of patients, adverse events, etc.
1.1 Related Work
Table 1: This table describes an overview of selected recent trial outcome prediction work. All previous work relies on industry sponsors to obtain labels, and most are not easily accessible. We are the first to aggregate publicly available data sources on a large number of trials.
Dataset Data Sponsor Subset # Trials Labeling Method
Pubicly
Available
Lo et al. [24] TrialTrove, Pharmaprojects Phase II, III 19,136 Human Expert
Feijoo et al. [7] Biomedtracker Industry Phase II, III 6,417 Manual linking
Aliper et al. [1] Insilico Medicine Phase II 55,653 Biomedical KG, Trial stats
Willigers et al. [43] Astrazeneca AstraZeneca Phase III 57 Human Expert
TOP [11] IQVIA Small Molecule Drugs 17,538 Human Expert
CTO (ours) Publicly Collected All 479,761
Publications, News, Trial linking,
Stocks, Trial stats, Etc
Predicting clinical trial outcomes is often led by industries with the capacity for extensive data curation. Informa’s TrialTrove [39] is widely used, containing around 20,000 trials [24]. AstraZeneca has developed structured feedback forms to improve Phase 3 trial success annotations [43]. Feijoo et al. [7] demonstrated that using Random Forest on Biomedtracker, a proprietary dataset aggregating company reports, results in strong outcome prediction performance.
Previous studies have tackled clinical trial outcome prediction using various methods. Early work employed statistical analysis [26] and ML models on limited private data sets (<500 samples) [6]. In drug toxicity prediction, Gayvert et al. [12] used Random Forest to predict outcomes based on chemical properties, while Artemov et al. [3] used Multilayer Perceptrons for Phase I/II trials. Lo et al. [24] applied KNN imputation and Random Forest on features from Pharmaprojects and TrialTrove. Additionally, Phase 2 to Phase 3 prediction was explored by Qi et al. [29] through clinical trials and by Aliper et al. [1] using experts, GPT-3.5, and a biomedical knowledge graph.
Fu et al. [11] released the first publicly available clinical trial outcome dataset based on manual curation. Train and test data splits were selected as all trials completed before 2014 and afterward respectively. After preprocessing and cleaning, the final number of trials in train, validation, and test datasets was reduced to 17,538 $\rightarrow$ 12,465 trials, which was only around 4% of all available trials at the time of publication 333and around 3% as of June 2024. Table 1 summarizes recent trial outcome work.
2 CTO Overview
Our main methodology is outlined in Figure 1. We overview our primary sources of outcome predictors–LLM prediction on Pubmed Abstracts, Trial Linkage, News Headlines, Stock Price, and finally, Trial Metrics computed from clincialtrials.gov. These metrics are computed independently and are aggregated via weakly supervised label aggregation.
2.1 LLM Predictions on PubMed Abstracts
@@
PubMed abstracts have been automatically linked to trials by the Clinical Trials Transformation Initiative (CTTI)[5, 40, 2] as well as other efforts [17]. First, We extracted all PubMed abstracts for each trial through the NCBI API  [^4] https://ncbi.nlm.nih.gov/ and the statistics of the extracted abstracts are given in the supplementary E.3. These abstracts can be categorized into 1) Background, 2) Derived, and 3) Results. Since we are interested in clinical trial outcomes, we only utilized abstracts in the Derived and Results categories. As many trials had multiple abstracts, we selected the top 2 abstracts based on their title similarity to the trial’s official title to provide the most relevant information.
Given these abstracts, we prompted the ‘gpt-3.5-turbo’ model to summarize important trial-related statistical tests and predict the outcome. Additionally, we prompted the LLM to generate QA pairs about the trial, which are provided as a supplement  C.3 to our benchmark. The prompts are provided in the supplementary G.
2.2 Trial Linkage
The journey of a drug from discovery to FDA approval involves several stages, beginning with Phase 1 trials to assess safety and dosage. Subsequent Phase 2 and 3 trials evaluate efficacy and compare the new drug to existing therapies. Upon completing Phase 3, a drug may be submitted for FDA approval. A key limitation of the CITI dataset is the lack of connectivity between trial phases, which could significantly enhance the ability to analyze trial progression and outcomes based on advancement to subsequent phases. Moreover, linking trials across phases is not straightforward due to challenges, including unstructured data, inconsistent reporting standards, missing information, data noise, and discrepancies in intervention details across phases. Despite these challenges, linking trials can be invaluable, particularly as a source of weak labels in clinical trial outcome prediction tasks. This section presents our novel trial-linking algorithm, which, to our knowledge, is the first attempt to systematically connect different phases of clinical trials. The trial linkage extraction process consists of two primary steps: 1) Linking trials across different phases, as illustrated in Figure 2B, and 2) Matching phase 3 trials with FDA approvals, as shown in Figure 2C.
Linking of Clinical Trials Across Phases
The progression of clinical trials through phases are not all strictly sequential from Phase 1 to Phase 3, as some studies may combine multiple phases. Many trials were categorized as ‘Not Applicable’ or were missing phase information entirely; these were excluded from our analysis. We created a phase connection map, illustrated in Figure 2A, that covers all phase categories present in the dataset. Our linking algorithm begins with the later phases and traces back to earlier phases. This approach is based on the assumption that a trial in a subsequent phase implies the success and existence of a corresponding trial in a preceding phase. Phase 3 trials are considered a success if a Phase 4 trial is found or if it is linked to an FDA new drug application.
The trial linkage process consists of three main steps: 1) Selection of the search space, 2) Retrieval of the top K most similar past trials, and 3) Prediction of linkage. In Figure 2B, we illustrate an example of linking a Phase 4 trial to its preceding phases. For a given trial $x$, the objective is to identify its predecessor among trials in its directly linked earlier phases. For instance, the directly linked earlier phases of Phase 4 are Phase 3 and Phase 2&\&&3.
Figure 2: Overview of trial linkage algorithm. A) Illustrates the phase connection map, covering all the phase categories present in the dataset. B) Linking trials across phases based on completion dates and intervention types, disease, and other trial data. We 1) constraint the search space, then 2) retrieve top-K most similar past trials using embeddings similarity. Finally, 3) a cross-encoder re-ranking strategy predicts linkages by scoring likely pairs. C) Matching FDA-approved drugs from FDA orange book to Phase 3 and Phase 2 &\&& 3 drug trials.
1. Search space selection: We filter the trials based on their completion dates relative to the start date, ensuring that any linkage candidate must have concluded before the start of $x$. Furthermore, we also consider intervention types such as ‘Drug’, ‘Biological’, ‘Device’, etc.
2. Retrieve top-K: From the filtered search space $\mathbf{Z}$, we retrieve top-32 most similar past trials to $x$. We extract key features and encode them into dense embedding using PubMedBERT [14] to represent both $x$ and trials in the search space ($z^{i}\in Z$) as follows: $x=\{x_{\text{I}},x_{\text{C}},x_{\text{T}},x_{\text{S}},x_{\text{E}}\}$, $z^{i}=\{z^{i}_{\text{I}},z^{i}_{\text{C}},z^{i}_{\text{T}},z^{i}_{\text{S}},z^{i}_{\text{E}}\}$. Where the subscript $I$ denotes intervention or drug, $C$ denotes condition or targeted disease, $T$ denotes official trial title, $S$ denotes trial summary, and $E$ denotes eligibility criteria. We calculate similarity as: $\text{similarity}(x^{i},z^{i})=\sum_{j\in\mathbf{F}}\frac{x^{i}\cdot z_{j}^{i}}{\|x^{i}\|\|z_{j}^{i}\|}$, where $\mathbf{F}=$I,C,T,S,E}.
We excluded the lead sponsor as a feature since the sponsor often changes depending on funding and performs worse empirically (Appendix F).
[…]
We apply this process for all trials in Phase 4, Phase 3, Phase 2 & 3, and Phase 2 to extract the trial linkages. To extract the outcome labels, we start with trials in the earlier phases and label them based on the existence of linked trials in the subsequent phases. However, trials in Phase 3 and Phase 4 have some exceptions to this process. For Phase 4 trials, there are no following trial phases, so we exclude them from the extracted weak labels. As for Phase 3 trials, they can be successful even without the existence of a subsequent Phase 4 trial. This highlights the importance of matching Phase 3 trials with their corresponding FDA approvals if they exist.
Matching Phase 3 trials with FDA approvals
After establishing connections across different phases of clinical trials, we focus on matching the Phase 3 trials to drug approvals to obtain their outcome labels. We utilize the FDA Orange Book [^5] https://fda.gov/drugs/drug-approvals-and-databases/orange-book-data-files version as of April 2024. Specifically, we use the approval date and drug name provided in the ‘product.txt’ file, as the other files do not contain the relevant information required for the matching process. In this process, we only consider drug-related trials in Phase 3 and Phase 2 & 3 since the Orange Book solely comprises FDA-approved drugs. For a given FDA-approved drug, we first filter Phase 3 and Phase 2 & 3 trials based on the approval date and intervention generic name, retaining trials completed between 2 years and 2 months prior to the approval date to align with the FDA approval process timeline. Similar to trial linkage prediction, we provide the drug’s generic name and the trial’s intervention generic name as input pairs to a cross-encoder, predicting their similarity. We select the top 5 trials based on cross-encoder scores and match the FDA approval to the trial with the completion date closest to the approval date. We then update the previously extracted outcomes for Phase 3 and Phase 2 & 3 trials, labeling them as successful if matched to an FDA approval or having a linked Phase 4 trial.
2.3 News Headlines
News headlines were obtained via the following steps: 1. Web Scraping: We sent requests to Google News for headlines regarding the top 1000 industry sponsors, which accounted for around 80% of the industry-sponsored trials (27,720 trials). Due to rate limitations, we limit our requests to a rate of around 1 query every 3-5 seconds. To obtain the widest range of news, we search each sponsor’s name and obtain up to 100 articles for every month, starting from the sponsor’s earliest clinical trial to the current day. We retrieved a total of 1,115,017 news articles. 2. News Sentiment Classification: We utilize FinBERT [45] to obtain financial news sentiment (‘Positive’ or ‘Negative’ with a confidence score between 0 and 1) for every headline. We drop the ‘Neutral’ sentiment as it is irrelevant for our task. 3. News / Trial Matching: Similar to Trial Linkage, we adopt a similar strategy of filtering using a top-K retriever and reranking using a cross-encoder. Trials are encoded with $x=\{x_{\text{I}},x_{\text{C}},x_{\text{T}},x_{\text{S}}\}\rightarrow z^{i}$, and headlines are encoded as $h^{i}$. We use PubMedBERT to encode both trials and headlines and follow the steps in trial linkage to obtain relevancy scores. We consider all headline sentiments with relevance scores larger than the mean score and take the mode of the sentiment predictions.
2.4 Stock price
@@
The stock price of a pharmaceutical or biotech company often reflects market expectations. If investors expect positive results, the stock may rise in anticipation of the trial’s completion. Conversely, if expectations are low or if previous trials have been unsuccessful, the stock may not perform as well. We utilized Yahoo finance [^6] https://pypi.org/project/yfinance/ to collect historical stock market data for companies that have publicly available tickers for completed trials. By averaging the stock prices over a specified time frame, the moving average reduces the noise caused by random, short-term price movements, making it easier to identify the underlying trend [19, 36]. A 5-day simple moving average (SMA) of a stock’s price is calculated by taking the average of the closing prices for the last 5 trading days. As shorter periods reveal shorter-term trends, we selected a 5-day SMA to capture the immediate short-term impact of the completion of clinical trials. A positive slope indicates an uptrend, while a negative slope indicates a downtrend in the SMA line [36]. The absolute value of the slope represents the steepness of the trend. We calculated the slope for a 7-day window starting at a clinical trial’s ‘completion date’.
2.5 Clinical Trial Metrics
We utilize the CTTI dataset to obtain preprocessed tables of trial details (e.g. eligibility criteria, statistics, linked references [^7] https://aact.ctti-clinicaltrials.org) [5, 40, 2]. Specifically, we utilize the following information. (1) Whether results were reported, (2) The number of sponsors, (3) The number of patients, (4) The patient dropout rate, (5) The number of sites or locations for the trial, (6) Whether the P-value $<$ 0.05, (7) The date at which the trial was last updated vs its completion date, (8) Number of deaths, (9) Number of serious adverse events, (10) Number of any adverse events, (11) The status of the trial e.g. terminated/withdrawn/completed/etc, and finally, (12) The number of amendments made to the trial page. Please see Appendix D.2 for in-depth discussions for a discussion on how trial outcomes are linked to these metrics.
Each of these metrics is treated as a weakly supervised Labeling Function (LF). For most of these metrics, we consider the "good" outcome as having greater or less than the median of that metric. For example, in the Serious Adverse Events LF, we predict "1" if a trial’s number of serious adverse events is less than the overall median number of serious adverse events. Otherwise, we predict "0". While choosing a highly specific threshold could be better than choosing the median, we note that LFs in the data programming framework (as we describe in the next section) do not have to be perfect, only better than random for data programming to work well.
2.6 Weakly Supervised Label Aggregation
We integrate multiple sources of weak supervision to select highly confident predictions from the previous step of relation prompting. Data programming is a framework designed to create denoised pseudo-labels from various weak supervision sources via labeling functions and matrix completion [32, 31].
A labeling function (LF) is a noisy heuristic that either assigns labels to unlabeled data or abstains from making a prediction. For example, f(text) = return SPAM if "HTTP" in-text else ABSTAIN is a labeling function used for spam detection. The main idea is that LFs that agree more with other LFs should be weighted higher, and given more weight in the final label prediction. Further details are introduced in Appendix D.1. To demonstrate the superiority of this approach, we also compare with Majority Vote baseline, where it simply takes the mode prediction of non-abstaining label prediction on any given data point. Finally, since we have train/test splits on TOP data, we can utilize these ground truth labels to enable a classification-like approach to weakly supervised label aggregation i.e. training a classifier to predict true labels from weakly supervised output. We use the Random Forest classifier for this task and ensure that no data leakage occurs by only training on the training set of TOP.
In total, we have more than 450k trials with automatically labeled outcomes. However, to enable comparison with the supervised labels in TOP, we analyze our trial labels on the interventional small-drug trials (However, we still release the full set of predicted trials).
3 Results
3.1 Agreement with Human Annotations

### v2

@@ Abstract
Background The global cost of drug discovery and development exceeds \$200 billion annually, with clinical trial outcomes playing a critical role in the regulatory approval of new drugs and impacting patient outcomes. Despite their significance, large-scale, high-quality clinical trial outcome data are not readily available to the public, limiting advances in trial outcome predictive modeling.
Methods We introduce the Clinical Trial Outcome (CTO) knowledge base, a fully reproducible, large-scale (around 125K drug and biologics trials), open-source of clinical trial information including large language model (LLM) interpretations of publications, matched trials over phases, sentiment analysis from news, stock prices of trial sponsors, and other trial-related metrics. From this knowledge base, we additionally performed manual annotation of a set of recent clinical trials from 2020-2024.
Results We evaluated the quality of our knowledge base by generating high-quality trial outcome labels that demonstrate strong agreement with previously published expert annotations, achieving an F1 score of 94 for Phase 3 trials and 91 across all phases. Additionally, we benchmarked a suite of standard machine learning models on our manually annotated set, highlighting the distribution shift of recent trials and the need for continuously updated labeling methods.
Conclusions By analyzing CTO’s performance on recent trials, we showed a need for recent, high-quality trial outcome labels. We release our knowledge base and labels to the public at https://chufangao.github.io/CTOD, which will also be regularly updated to support ongoing research in clinical trial outcomes, offering insights that could optimize the drug development process.
Introduction
@@
A clinical trial is an indispensable step toward developing a new drug, involving human participants to test the drug’s efficacy and safety for treating target diseases. In 2022, drug discovery and development spending reached 244 billion dollars globally [1]. Among these, the clinical trial market reached \$44.3 billion in 2020 and is expected to grow to \$69.3 billion by 2028 [2]. Low efficacy, safety issues, and poor trial protocol design can lead to trial failures [3, 4, 5]. Eroom’s Law111reverse of “Moore’s Law” shows that the number of new FDA-approved drugs per billion US dollars of R&D spending has halved approximately every nine years since 1950, even with inflation adjustment [6]. Despite this decline in approval efficiency, the volume of drug and biologics trials continues to grow each year. However, obtaining high-quality, human-labeled trial data remains expensive and largely inaccessible due to the limited public availability of labels. Given these challenges, predicting trial outcomes in silico—using computational methods—could significantly enhance drug development efficiency.
The FDA does not release the clinical trial ID (NCTID) in documents of approved drug applications. Other public data sources, such as the ClinicalTrials.gov database with more than 500,000 historical trials [7, 8, 9], provide vital information for identifying trial recruitment and design information but does not provide a concrete label for trial outcome directly. The absence of a centralized, easily accessible knowledge base that consolidates clinical trial outcomes, drug approvals, and intervention phases poses a significant challenge for researchers and drug developers [10]. This fragmented information landscape hinders the development of accurate predictive models and can slow down the drug discovery and development process. Furthermore, the number of clinical trials conducted each year continues to rise, with their data dynamically evolving under the influence of various external factors, such as technological innovations and global events like the COVID-19 pandemic. This ever-changing landscape underscores the need for updated trial outcome labels to develop robust and high-performing clinical trial outcome prediction models.
@@
Trial Outcome Definition: Clinical trial outcomes are multifaceted and have diverse implications. These outcomes can involve meeting the primary endpoint as defined in the study, advancing to the next phase of the trial, obtaining regulatory approval, impacting the financial outcome for the sponsor (either positively or negatively), and influencing patient outcomes such as adverse effects and trial dropouts. Our paper follows the previous conventions [11, 12, 3, 13, 14, 15] and defines the trial outcome as a binary indicator, showing whether the trial achieves its primary endpoints and can progress to the next stage of drug development. For Phase 1 and 2 trials, success may mean moving to the next phase, such as from Phase 1 to Phase 2, and from Phase 2 to Phase 3. In Phase 3, success is measured by regulatory approval.
@@
Recently, initial efforts have been made to forecast specific aspects of clinical trials to enhance their outcomes. These efforts include employing electroencephalographic (EEG) measures to predict the effects of antidepressant treatments [16], optimizing drug toxicity predictions based on drug and target properties [17], and using phase II trial results to anticipate phase III trial outcomes [18]. Additionally, there is a growing interest in developing comprehensive methods for predicting trial outcomes. For example, predicting drug approvals for 15 disease groups by analyzing drug and clinical trial features using classical machine learning techniques [14], using multimodal drug structure and text information to predict outcomes based on a supervised set of data [11], an algorithm that computes the probability of technical success via asking experts a standardized questionnaire [19], and multimodal trial outcome prediction via omics, text, clinical trial design, and small molecule properties [15].
Despite these efforts, several limitations still impede the utility of existing trial outcome prediction models. Namely–the lack of transparency in the clinical trial labeling process. Fu et al.’s TOP [11] is one example of a publicly available expert-curated clinical trial dataset. However, it is also quite limited because it only contains around 12,000 human-labeled interventional small-molecule drug trials, most of which concluded prior to 2020. To date, we are unaware of any other large-scale, publicly available, open-source (fully reproducible) effort to compute trial outcome labels.
To address these limitations, we propose CTO, the first large-scale, publicly available, open-source, and fully reproducible dataset of clinical trial outcomes. We compiled a comprehensive trial knowledge base by integrating diverse data sources, including trial-related publications, news headlines, stock prices of trial sponsors, and other relevant trial metrics. From this knowledge base, we derive weakly supervised trial outcome labels using the CTO’s automated labeling framework, which aggregates labels from trial phase linkages, LLM interpretations on publications, sentiment analysis of news headlines, p-values, and trial status, among other indicators. We validate CTO ’s labeling accuracy against previously published trial outcomes, achieving an F1 score of 91 across all trial phases, closely aligned with human-annotated clinical trial outcomes. Additionally, aided by the CTO’s framework, we manually curated a set of approximately 10,000 recent trials completed between 2020 and 2024. To ensure reproducibility and keep pace with rapidly evolving trial data, we share monthly updates of our trial knowledge base and labels generated by CTO ’s automated labeling framework.
Figure 1: Overview of our CTO benchmark pipeline. (A) Trial selection criteria used to curate a targeted dataset from the CTTI, narrowing from an initial 513.4K trials to 125.8K trials in CTO to a final manually curated set of 11K trials. (B) Trial knowledge base creation pipeline, including extraction of linked PubMed abstracts, recent abstract searches, trial metrics, news articles, and stock prices. (C) Clinical trial outcome labeling process, integrating the CTO automated labeling framework assisted manual labeling to generate curated gold standard labels. (D) The manual curation process, our rule-based annotation criteria, followed by manual checking on challenging clinical trials.
Methods
Trial Selection Process
In our study, we focused primarily on clinical trials involving drugs and biologics. Starting with a dataset of approximately 513,485 clinical trials from the CTTI, we applied a series of selection criteria to create a refined subset as illustrated in Figure 1A. First, we selected only trials with interventions categorized as drugs or biologics, reducing the dataset to 209,124 trials. Next, we excluded active trials and those still recruiting to focus exclusively on completed studies, resulting in 145,464 trials. Subsequently, trials without phase information were removed, further narrowing the dataset to 125,840 trials, which represents the number of trials labeled using our CTO’s automated labeling framework. To construct a manually curated test set, we then selected trials completed after 2020, yielding a subset of approximately 30,900 recent trials. Importantly, we ensured that this set had no overlap with the TOP benchmark. Finally, we performed a sampling across different trial phases and conditions, producing a manually curated set of 11,012 trials.
Trial Knowledge Base Creation
Based on the CTTI data, we generate a trial knowledge base by aggregating and linking various data sources to the trials, as shown in Figure 1B. Our pipeline enables regular updates of the knowledge base given the updated CTTI data in a monthly frequency. The trial knowledge base extraction pipeline consists of four components: 1) PubMed abstracts extraction, 2) News extractions, 3) Trial metrics and 4) Stock price information extraction.
@@ Methods
Pubmed abstract extraction: PubMed abstracts have been automatically linked to trials by the Clinical Trials Transformation Initiative (CTTI)[20, 21, 22] as well as other efforts [23]. Despite these efforts, we found that many recent abstracts were not linked to their corresponding trials. First, we extracted all linked PubMed abstracts for each trial through the NCBI API  [^2] https://ncbi.nlm.nih.gov/, and the statistics of the linked abstracts are given in the Supplementary. These abstracts can be categorized into 1) Background, 2) Derived, and 3) Results. Since we are interested in clinical trial outcomes, we only utilized abstracts in the Derived and Results categories. To further enrich the knowledge base, we performed regular searches on NCBI using NCT IDs to find newly published abstracts, linking any relevant ones to their associated trials. The initial count of drug or biologic trials with extracted PubMed abstracts linked from CTTI data was 49,277. Our additional search and extraction process linked 10,232 new trials with their recent PubMed abstracts. Ultimately, our knowledge base comprises 41,228 Results-type abstracts, 65,327 Derived-type abstracts, and 20,512 newly linked abstracts.
News extractions: In our initial analysis, we identified news sources as a valuable resource for determining the outcome of clinical trials, especially since the majority of trials in the CTTI dataset lack reported results. We extracted news articles related to completed drug and biologic trials by web scraping and linked them to the trials in our knowledge base. We sent requests to Google News for all completed drug and biologic trials (excluding phase 4), which accounted for around 85K trials. Due to rate limitations, we used an external API (SerpAPI) to streamline this process. Our findings revealed that most trials had no corresponding news coverage; only 3,688 trials were associated with any news articles.
Trial metrics: We extracted trial metrics from CTTI[24] as well as mining clinicaltrials.gov. Specifically, we extracted a total of 9 numerical metrics–number of patients, sponsors, sites, amendments, patient drops, deaths, serious adverse events, all adverse events, and the number of days the trial has been last updated after completion. We also extracted 3 binary metrics–whether results were reported, the completion/terminated status of the trial, and finally, whether the trial had a significant p-value. Detailed descriptions of each are in the supplementary.
@@
Stock price information extraction: The stock price of a pharmaceutical or biotech company often reflects market expectations. If investors expect positive results, the stock may rise in anticipation of the trial’s completion. Conversely, if expectations are low or if previous trials have been unsuccessful, the stock may not perform as well. We utilized Yahoo finance [^3] https://pypi.org/project/yfinance/ to collect historical stock data for companies with publicly available tickers for completed trials. Details on the calculations involving stock price movements and their linkage to specific trials are provided in the supplementary.
CTO Automated Labeling Framework
We developed a clinical trial labeling pipeline that first assigns rule-based labels. First, trials with statuses indicating ‘terminated’, ‘withdrawn’, ‘suspended’, ‘no longer available’} were labeled as failures. Trials labeled as ’approved for marketing’ were marked as successes. For trials with ‘completed’ as status, we examined the reported results: if any reported p-value for a primary endpoint was $<$0.05, the trial was labeled as a success. Unlabeled trials are then automatically labeled. The main idea behind our automated labeling framework is to combine multiple sources of labels, which have varying levels of accuracy. We call these sources of labels “weakly supervised labeling functions” or LFs for short[25]. Following this, we apply a label model that leverages our sources to refine labels and produce denoised pseudo-labels. Each LF, such as detecting efficacy based on p-value thresholds, assigns labels or abstains when uncertain. Phase-specific thresholds are also applied to certain metrics, such as adverse events, to account for variations in participant numbers across trial phases.
From our trial knowledge base, we developed several sources of labels, including large language model (LLM) interpretations of PubMed abstracts, trial linkage across phases, sentiment analysis of news headlines, p-values, amendments, and other clinical trial metrics which were integrated into our clinical trial annotation pipeline as shown in Figure 1C. Further details on these labeling functions are given in the supplementary. Our pipeline is validated against several baselines, including a Majority Vote–which takes the majority vote over LFs, Data Programming–an unsupervised matrix completion approach that weighs LFs based on mutual agreement, and a supervised Random Forest model trained on LFs and a set of public labels (TOP).
Results
CTO Automated Labeling Achieves High Agreement with Previous Expert Labels